OpenStudy (anonymous):

if the area of a base of a rectangular solid is tripled, what is the percent increase in the volume?

OpenStudy (anonymous):

here is a trick to use. since it does not matter what the base is make it a number. say you have a cube with edge length 2 one and so the area of the base is 2^2=4 and the volume is 2^3=8. then you triple the area of the base to 12. now the side of edge of the cube has length \[\sqrt{12}=2\sqrt{3}\] and the volume is \[\sqrt{12}^3=(2\sqrt{3})^3=8\sqrt{3}^3=72\sqrt{3}\] you want the percent increase so divide \[\frac{72\sqrt{3}}{8}=9\sqrt{3}\]

OpenStudy (anonymous):

why you want to write this number as a percent is beyond me, but if you triple the are of the base the volume increases by a factor of \[9\sqrt{3}\] or \[9^{\frac{3}{2}}\]

OpenStudy (anonymous):

but the question is asking for the percent increase. it's either 200, 300, 600 or 900.

OpenStudy (anonymous):

900

OpenStudy (anonymous):

the answer key says 200 for some reason.

OpenStudy (anonymous):

then i guess i am wrong...

OpenStudy (anonymous):

no in fact i am right. take a cube with side edge 1, so base has side 1 and volume is 1. triple the area of the base. it is 3, the edge is \[\sqrt{3}\] and the volume is \[\sqrt{3}^3=9\sqrt{3}\] that it the increase, from 1 to \[9\sqrt{3}\]

OpenStudy (anonymous):

The question is asking about a rectangle, not a cube. So can anyone kindly explain the correct answer?